### v3

# Strategic Analysis of Fair Rank-Minimizing Mechanisms with Agent Refusal Option

Abstract
This study examines strategic issues in fair rank-minimizing mechanisms, which choose an assignment that minimizes the average rank of object types to which agents are assigned and satisfy a fairness property called equal treatment of equals. As one of these fair mechanisms, the uniform rank-minimizing mechanism is considered. We focus on the case where agents can refuse their assignment and obtain the outside option instead. Without the refusal option, truth-telling is not strategically dominated by any strategies if a fair rank-minimizing mechanism is used. However, if agents have the option and the uniform rank-minimizing mechanism is used, then a strategy called an outside option demotion strategy strategically dominates truth-telling. Moreover, we show that adopting this strategy may lead to inefficient assignments. To counter this, we propose a modification of the uniform rank-minimizing mechanism, though it may lead agents to strategically reduce the number of acceptable types.
JEL classification: C78; D47
Keywords: Rank-minimizing mechanism; Strategic manipulation; Random assignment; Outside option; Refusal option
1 Introduction
In this study, we discuss the strategic issues of fair rank-minimizing allocation mechanisms mainly when agents have the option to refuse their assignment determined by one of the mechanisms and obtain their outside option instead. There are several agents and types of object, including a null object type, which corresponds to the outside option. In the mechanism, each agent reveals one preference ranking, which may or may not be their true preference ranking. In rank-minimizing mechanisms, assignments are determined as to minimize the total (or average) revealed rank of the objects to which the agents are assigned.
The rank distributions of the assignments are seriously considered by many real world matching authorities. For example, Featherstone (2020) mentions that Teach for America considers the rank distribution when selecting its assignment. Therefore, the rank-minimizing mechanisms are focused on by several recent studies such as Featherstone (2020), Nikzad (2022), Ortega and Klein (2023) and Troyan (2024). Some of them highlight several advantages of the mechanism with regard to the efficiency of assignment. Featherstone (2020) shows that any rank-minimizing mechanism satisfies another property of efficiency called ordinal efficiency, first introduced by Bogomolnaia and Moulin (2001). Moreover, Ortega and Klein (2023) compare the result of a rank-minimizing mechanism and those of the DA mechanism and the TTC mechanism, and show that the result of a rank-minimizing mechanism improves upon those of others in several terms of efficiency.
In this study, we especially consider the rank-minimizing mechanisms that satisfy a fairness property called equal treatment of equals, which requires that if two agents reveal the same preference ranking, then they are assigned to each object type with the same probability. This fairness property is discussed by Bogomolnaia and Moulin (2001) and many other previous studies on various allocation problems (Thomson 2011). We consider a slightly stronger version of this property. We call a rank-minimizing mechanism satisfying equal treatment of equals a fair rank-minimizing mechanism. To consider these mechanisms, we adopt the probabilistic assignment framework. Among the mechanisms, we first consider a representative one called the uniform rank-minimizing mechanism. Since Nikzad (2022), Ortega and Klein (2023) and Troyan (2024) also mainly consider this mechanism, at least in the literature, this is the most common rank-minimizing mechanism.
We discuss the strategic manipulation incentives of agents when the assignment is determined by a fair rank-minimizing mechanism. Specifically, we consider the case where agents can refuse their assignment, and if they refuse it, then the outside option is assigned to them. First, as a benchmark case, we discuss the situation where agents cannot refuse their assignment. This case is discussed by many previous studies, such as those mentioned above. Featherstone (2020) shows that any rank-minimizing mechanism is not strategy-proof. On the other hand, Troyan (2024) shows that the uniform rank-minimizing mechanism is not obviously manipulable, which is a strategic property introduced by Troyan and Morrill (2020). Moreover, Nikzad (2022) shows that the uniform rank-minimizing mechanism is Bayesian incentive compatible in uniform markets, which he primarily discusses. Related to but logically independent to their results, we show that, in this case, truth-telling is not strategically dominated by any untrue preference under any fair rank-minimizing mechanism. These results imply that in this case, strategic manipulation may not be a critical problem of rank-minimizing mechanisms.
Next, we consider the case where agents can refuse their assignment and instead obtain the outside option. Featherstone (2020) firstly considers this situation, but focuses only on the case where each agent has limited information on both the preferences of other agents and the capacities of all object types except the outside option. In contrast, we assume that each agent has complete information on the capacities of all object types. For example, in a school choice market, information about all schools, including their capacities, is assumed to be disclosed to all students. Additionally, we do not need any assumptions about the information structure of agents regarding the preferences and rationality of other agents because we focus only on strategic dominance relations.
We consider an outside option demotion strategy, which involves revealing a ranking where the rank of the outside option is placed at the bottom, without changing the order of the truly acceptable object types. We show that, if the uniform rank-minimizing mechanism is used and agents have the refusal option, then any outside option demotion strategy weakly strategically dominates truth-telling. Moreover, with some additional conditions, some outside option demotion strategy (strictly) strategically dominates truth-telling.
The reason is as follows: first, for some preference profiles of other agents, an agent succeeds in avoiding the outside option and getting a truly acceptable object type because assigning the outside option to them significantly increases the sum of ranks. Of course, due to this manipulation, the agent faces the risk of being assigned an unacceptable object type. However, in the case where the agents have the refusal option, the agent can refuse any of such types and obtain the outside option instead. Moreover, we show that if an agent adopts an outside option demotion strategy and the uniform rank-minimizing mechanism is used, an inefficient assignment may be realized. This is because, in that case, an unacceptable type may be assigned to the agent who adopts the strategy, and then the agent refuses it, even though the type is desired by other agents. Therefore, in this case, strategic manipulation is a critical problem for rank-minimizing mechanisms.
Featherstone (2020) shows that revealing the full extension, which is one of the outside option demotion strategies, is a weakly better strategy than any other strategies in the case where the agents have little information and the refusal option. The implication of Featherstone (2020) is that rank-efficient mechanisms have strategic issues if the agents have little information and the outside option is somewhat valued by some agents. On the other hand, by our result, this implication is robust for the uniform rank-minimizing mechanism even without the low information assumption. Moreover, we show that revealing the full extension may be strategically dominated by another outside option demotion strategy because, contrary to our model, Featherstone (2020) assumes that the agents have no information on the capacities of object types other than their outside option.
Due to the negative result on the uniform rank-minimizing mechanism introduced above, we modify it so that truth-telling is not strategically dominated by any other strategy. We call this mechanism the modified uniform rank-minimizing mechanism, which is shown to be fair. That is, to ensure that any outside option demotion strategy can fail, in the mechanism, increasing acceptable object types can make the agent less likely to be assigned their acceptable types. However, using the modified uniform rank-minimizing mechanism may create another strategic issue: an agent may have an incentive to reduce the number of truly acceptable object types.
2 Model
We consider the model discussed by Kojima and Manea (2010) and Featherstone (2020). Let $\mathcal{A}$ and $\mathcal{O}$ be finite sets of agents and object types respectively. We assume $\left|\mathcal{A}\right|\geq 2$ and $\left|\mathcal{O}\right|\geq 2$. Each agent must be allocated one copy of an object whose type is an element of a set denoted by $\mathcal{O}$. Let $q_{o}\geq 1$ be the capacity of an object type $o\in\mathcal{O}$. We assume that there is a null object type denoted by $\varnothing\in\mathcal{O}$ satisfying $q_{\varnothing}\geq\left|\mathcal{A}\right|$. That is, the null object type can be assigned to all agents simultaneously. The null object type $\varnothing\in\mathcal{O}$ can be interpreted as the (best) outside option of an agent. In Section 4, we consider the situation where the null object type has the other characteristic. To avoid trivial cases, we assume $q_{o}\in\left[1,\left|\mathcal{A}\right|\right)$ for all $o\in\mathcal{O\setminus}\left\{\varnothing\right\}$.
Let $\mathcal{P}$ be the set of all possible strict rankings (linear orders) of the elements in $\mathcal{O}$. Let $p_{a}=\left(p_{a}^{1},\cdots,p_{a}^{\left|\mathcal{O}\right|}\right)\in\mathcal{P}^{\left|\mathcal{A}\right|}$ be a preference order revealed by $a$, where $p_{a}^{k}=o$ means that type $o$ object is ranked $k$th best one under $p_{a}$. For notational convenience, we let $R:\mathcal{O\times P}\rightarrow\left\{1,\cdots,\left|\mathcal{O}\right|\right\}$ be a rank function; that is, $p_{a}^{k}=o$ if and only if $R\left(o,p_{a}\right)=k$. Moreover, if $R\left(o,p_{a}\right)<R\left(\varnothing,p_{a}\right)$ (resp. $R\left(o,p_{a}\right)>R\left(\varnothing,p_{a}\right)$), then $o$ is said to be acceptable (resp. unacceptable) for $p_{a}$. Thus, $R\left(\varnothing,p_{a}\right)-1$ represents the number of acceptable object types for $p_{a}$. Let $r_{a}\in\mathcal{P}$ be the true preference order of agent $a$. Especially, if $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, then $o$ is said to be truly acceptable for $a$.
An assignment is represented as lotteries over types. Let $x_{ao}$ be the probability that agent $a$ assigns (a copy of) type $o$ object. Moreover, let $x$ be a matrix whose $\left(a,o\right)$ entry is equal to $x_{ao}$. We say that a matrix $x$ is an assignment if
$x_{ao}\in\left[0,1\right]\text{, }\sum\limits_{o^{\prime}\in\mathcal{O}}x_{ao^{\prime}}=1\text{,\ and }\sum\limits_{a^{\prime}\in\mathcal{A}}x_{a^{\prime}o}\leq q_{o},$
for all $a\in\mathcal{A}$ and all $o\in\mathcal{O}$. Let $\mathcal{X}$ be the set of all possible assignments. Specifically, let $y\in\mathcal{X}$ be a deterministic assignment if $y_{ao}$ is either 00 or $1$ for all $\left(a,o\right)\in\mathcal{A}\times\mathcal{O}$. Let $\mathcal{Y}\subseteq\mathcal{X}$ be a set of all possible deterministic assignments.
The following result is due to Kojima and Manea (2010), which is a generalization of the Birkhoff–von Neumann theorem.
Remark 1
Every assignment can be written as a convex combination of deterministic assignments.
We consider the preference of agent $a$ with $p_{a}$ with regard to two assignments $x$ and $x^{\prime}$. If
$\sum\limits_{o^{\prime}:\text{ }R\left(o^{\prime},p_{a}\right)\leq k}x_{ao^{\prime}}\geq\sum\limits_{o^{\prime}:\text{ }R\left(o^{\prime},p_{a}\right)\leq k}x_{ao^{\prime}}^{\prime}$ (1)
for all $k=1,2,\cdots,\left|\mathcal{O}\right|$, then agent $a$ with $p_{a}$ is said to weakly prefer (in the sense of first-order stochastic dominance) $x$ to $x^{\prime}$. Moreover, $a$ is said to strictly prefers $x$ to $x^{\prime}$ if $a$ weakly prefers $x$ to $x^{\prime}$ and (1) is satisfied with inequality for some $k=1,2,\cdots,\left|\mathcal{O}\right|-1$.
Assignment $x$ weakly ordinally dominates assignment $x^{\prime}$ for $p$ if each agent $a$ with $p_{a}$ weakly prefers $x$ to $x^{\prime}$. If $x$ weakly ordinally dominates assignment $x^{\prime}$ for $p$ and (1) is satisfied with inequality for at least one agent–rank pair, then $x$ ordinally dominates assignment $x^{\prime}$ for $p$. An assignment that is not ordinally dominated by any other assignments is said to be ordinally efficient for $p$.
Assignment $x$ is wasteful for $p$ if there are $o,o^{\prime}\in\mathcal{O}$ and $a\in\mathcal{A}$ such that $\sum\nolimits_{a^{\prime}\in\mathcal{A}}x_{a^{\prime}o}<q_{o}$, $x_{ao^{\prime}}>0$ and $R\left(o,p_{a}\right)<R\left(o^{\prime},p_{a}\right)$. Any wasteful assignment for $p$ must not be ordinally efficient. [^1] Moreover, a wasteful assignment for $p$ must not be efficient in a weaker efficiency concept such as ex-post efficiency, which is discussed several previous studies such as Bogomolnaia and Moulin (2001). Thus, not being wasteful is a very weak requirement for efficiency.
For an assignment $x\in\mathcal{X}$ and a preference profile $p=\left(p_{a}\right)_{a\in\mathcal{A}}\in\mathcal{P}^{\left|\mathcal{A}\right|}$, let $RV\left(x,p\right)$ be such that
$RV\left(x,p\right)=\sum\limits_{a\in\mathcal{A}}\sum\limits_{o\in\mathcal{O}}R\left(o,p_{a}\right)\times x_{ao},$
which is the sum of expected rank. An assignment $x^{\ast}\in\mathcal{X}$ is rank efficient for $p$ if $RV\left(x^{\ast},p\right)\leq RV\left(x,p\right)$ for all $x\in\mathcal{X}$. Let $\mathcal{X}^{\ast}\left(p\right)$ be the set of rank efficient assignments for $p$.
The following result is shown by Featherstone (2020).
Remark 2
Any rank efficient assignment for $p$ is ordinally efficient for $p$.
In the subsequent section, we focus on the mechanisms whose input is $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ and output is rank efficient for any $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$. Before that we introduce some technical results.
Lemma 1
Fix any $x^{\ast}\in\mathcal{X\setminus Y}$. Let $y^{1},\cdots,y^{I}\in\mathcal{Y}$ with $I\geq 2$ be such that
$x^{\ast}=\sum\limits_{i=1}^{I}\sigma^{i}y^{i}\in\mathcal{X}\text{ where }\sum\limits_{i=1}^{I}\sigma^{i}=1$
and $\sigma^{i}\in\left[0,1\right]$ for all $i=1,\cdots,I$. If $x^{\ast}$ is rank efficient, then $y^{i}$ is also rank efficient for all $i=1,\cdots,I$.
We provide the proofs of our results in the Appendix.
By Lemma 1, we immediately have the following result.
Corollary 1
There must exist at least one deterministic rank efficient assignment.
Nikzad (2022) and Troyan (2024) consider a deterministic rank efficient assignment; that is, they focus on a deterministic assignment $y^{\ast}\in\mathcal{Y}$ such that $RV\left(y^{\ast},p\right)\leq RV\left(y,p\right)$ for all $y\in\mathcal{Y}$. However, by Corollary 1, such $y^{\ast}$ must also be rank efficient in our definition; that is, $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for all $x\in\mathcal{X}$. Moreover, this result allows us to show that the mechanism introduced later is well-defined.
Finally, we introduce another technical result that is used to show our main results introduced later. Let
$\bar{k}\left(p_{a}\right)=\min\left\{k\in\left\{1,\cdots,\left|\mathcal{O}\right|\right\}\text{ }\left|\text{ }\sum\limits_{\bar{o}:R(\bar{o},p_{a})\leq k}q_{\bar{o}}\geq\left|\mathcal{A}\right|\right.\right\},$
which is well-defined because $\sum\nolimits_{\bar{o}\in\mathcal{O}}q_{\bar{o}}>\left|\mathcal{A}\right|$. This means that for a given $p_{a}$, there are sufficient number of copies of object types that ranked higher than or equal to $\bar{k}\left(p_{a}\right)$. Trivially, if $p_{a}^{k}=\varnothing$, then $\bar{k}\left(p_{a}\right)\leq k$. We have the following result.
Lemma 2
Let $x\in\mathcal{X}$ be an assignment that is not wasteful for $p$. Then, $x_{ao}=0$ for all $o$ such that $R(o,p_{a})>\bar{k}\left(p_{a}\right)$.
Then, if an assignment is not wasteful, then $a$ must be assigned to an object type that ranked higher than or equals to $\bar{k}\left(p_{a}\right)$. Typically, in that case, the object types that ranked lower than $\varnothing$ in the preference ranking of agent $a$ are never assigned to $a$, because $\varnothing$ is the object with a sufficient number of copies.
3 Rank-minimizing Mechanism
In this section, we consider the following game. First, all agents simultaneously reveal a preference order in $\mathcal{P}$, which may or may not be their true preference order $r_{a}\in\mathcal{P}$. Let $p_{a}\in\mathcal{P}$ be the preference order revealed by $a\in\mathcal{A}$ and $p=\left(p_{a}\right)_{a\in\mathcal{A}}\in\mathcal{P}^{\left|\mathcal{A}\right|}$. Second, based on $p$, an assignment is determined via a mechanism. We assume that all agents know the capacities of all object types $\left(q_{o}\right)_{o\in\mathcal{O}}$ and the mechanism used to determine the assignment, when they reveal a preference order. Note that in this section, the null object type is just that with sufficient number of copies.
A mechanism used in the second stage is represented by $f:\mathcal{P}^{\left|\mathcal{A}\right|}\rightarrow\mathcal{X}$. First, we consider an efficiency property on mechanisms. We call $f$ a rank-minimizing mechanism if $f(p)$ is rank efficient for all $p$.
Second, we introduce a fairness property. First, we introduce a well-known one. A mechanism $f$ is said to satisfy weak equal treatment of equals (hereafter weak ETE) if $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfying $p_{a}=p_{a^{\prime}}$ implies $\left(f\left(p\right)\right)_{ao}=\left(f\left(p\right)\right)_{a^{\prime}o}$ for all $o\in\mathcal{O}$. This requires that two agents revealing the same preference must have the same assignment in the ex-ante sense. This property is called just “equal treatment of equals” by many previous studies such as Bogomolnaia and Moulin (2001) and Troyan (2024). In this study, we mainly use a slightly stronger one. A mechanism $f$ is said to satisfy equal treatment of equals (hereafter ETE) if $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfying $p_{a}^{k^{\prime}}=p_{a^{\prime}}^{k^{\prime}}$ for all $1,\cdots,\bar{k}\left(p_{a}\right)$ implies $\left(f\left(p\right)\right)_{ao}=\left(f\left(p\right)\right)_{a^{\prime}o}$ for all $o\in\mathcal{O}$ [^2] Han (2024) also considers a similar property, called symmetricity, which is slightly stronger than equal treatment of equals. Unlike in our model, they take into account the priorities of objects and define agents as equal only if they are equal with respect to the priorities of all objects.
By Lemma 2, as long as the assignment is not wasteful, any type that is ranked lower than $\bar{k}\left(p_{a}\right)$th in the preference ranking revealed by $a$ will not be assigned to $a$, because there are sufficient copies of types ranked higher than or equal to $\bar{k}\left(p_{a}\right)$th. Thus, if two individuals reveal rankings that are identical for positions at or above $\bar{k}\left(p_{a}\right)$th rank, those rankings are considered essentially the same, and they should receive the same assignment in an ex-ante sense.
We focus on the efficient and fair mechanisms; that is, we call $f$ a fair rank-minimizing mechanism if $f(p)$ is rank efficient for all $p$ and satisfies ETE. The following mechanism is one of them.
Let $\mathcal{Y}^{\ast}\left(p\right)\subseteq\mathcal{Y}$ be the set of deterministic rank efficient assignments; that is, $y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right),$ $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$. By Corollary 1, $\mathcal{Y}^{\ast}\left(p\right)\neq\emptyset$. Let
$f^{U}\left(p\right)=\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}\text{ for all }p\in\mathcal{P}^{\left|\mathcal{A}\right|}$
be the uniform rank-minimizing mechanism. By (7),
$RV\left(f^{U}\left(p\right),p\right)=\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{1}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}RV\left(y^{\ast},p\right).$
Thus, $RV\left(f^{U}\left(p\right),p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$; that is, $f^{U}\left(p\right)$ must be rank efficient for $p$. Since Nikzad (2022), Ortega and Klein (2023) and Troyan (2024) also discuss the uniform rank-minimizing mechanism, this is one of the most common rank-minimizing mechanisms in the literature.
We have the following result.
Proposition 1
The uniform rank-minimizing mechanism satisfies ETE.
Therefore, the uniform rank-minimizing mechanism is a fair rank-minimizing one.
We discuss strategic problems of fair rank-minimizing mechanisms. We call the strategy of revealing their true preference truth-telling. Featherstone (2020) shows that any rank-minimizing mechanism does not satisfy strategy-proofness. Therefore, in this paper, we consider a weaker strategic property.
We discuss whether truth-telling ($r_{a}$) can be strategically dominated by an untrue preference $p_{a}\neq r_{a}$. Let $k$ be the smallest integer such that $p_{ak}\neq r_{ak}$. By Lemma 2, if $k>\bar{k}\left(p_{a}\right)$, then $p_{a}$ and $r_{a}$ are essentially equivalent under a fair rank-minimizing mechanism, because any result of a rank-minimizing mechanism is not wasteful. Thus, we focus only on the case where $k\leq\bar{k}\left(p_{a}\right)$ defined above; that is,
$\sum\limits_{\bar{o}:R(\bar{o},p_{a})<k}q_{\bar{o}}\left(=\sum\limits_{\bar{o}:R(\bar{o},r_{a})<k}q_{\bar{o}}\right)<\left|\mathcal{A}\right|$
is satisfied for that $k$. In other word, we consider the case where the number of object types that are more preferable than the $k-1$th best one in $p_{a}$ is insufficient to distribute to all agents. In this case, for some $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and some fair rank-minimizing mechanism, $a$ is assigned to an object type that is less preferable or equal to the $k$th best one.
Lemma 3
Suppose that $f$ is a fair rank-minimizing mechanism. Let $p_{a}\neq r_{a}$, and $k$ be the smallest integer such that $p_{a}^{k}\neq r_{a}^{k}$. If $k\leq\bar{k}\left(p_{a}\right)$, then there is $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that agent $a$ (with $r_{a}$) strictly prefers $f\left(r_{a},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$.
By this lemma, we have the following result.
Proposition 2
Suppose that a fair rank-minimizing mechanism is used. For any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ such that revealing $p_{a}$ strategically dominates truth-telling.
This result implies that any fair rank-minimizing mechanism satisfies a weak strategic characteristic; that is, for each untrue preference $p_{a}$, either $\left(f\left(r_{a},p_{-a}\right)\right)_{ao}=\left(f\left(p_{a},p_{-a}\right)\right)_{ao}$ for all $o\in\mathcal{O}$ and all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ or there must exist $p_{-a}$ such that $a$ is worse off by revealing it instead of $r_{a}$. Thus, if a fair rank-minimizing mechanism is used, any (essential) strategic manipulation by an agent carries some level of risk. This characteristic is quite weak, but in the subsequent section, we show that this is not satisfied when the agents can refuse their assignment and instead obtain the null object type.
Note that Proposition 2 holds for any fair rank-minimizing mechanisms but cannot be generalized to the rank-minimizing mechanisms satisfying weak ETE. We show this by introducing the following example.
Example 1
Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$, $\mathcal{O=}\left\{o_{1},o_{2},o_{3},\varnothing\right\}$ and $q_{o_{1}}=1$. Let $\psi=\left(o_{1},o_{2},\varnothing,o_{3}\right)$ and $\psi^{\prime}=\left(o_{1},o_{2},o_{3},\varnothing\right)$. A rank-minimizing mechanism $f$ is such that
$\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f\left(\psi,\psi^{\prime},\psi\right)\right)_{a_{2}o_{1}}=\left(f\left(\psi,\psi,\psi^{\prime}\right)\right)_{a_{3}o_{1}}=0,$ (2)
$f\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ $=$ $\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}$
if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ is neither $\left(\psi^{\prime},\psi,\psi\right)$ nor $\left(\psi,\psi^{\prime},\psi\right)$ nor $\left(\psi,\psi,\psi^{\prime}\right)$. That is, if the best type and the second best type of the three agents are respectively $o_{1}$ and $o_{2}$ and the third best type of the two of them is $\varnothing$, then the agent revealing $o_{3}$ as the third best one obtains $o_{1}$ with probability 00. Trivially, $f$ satisfies weak ETE. However, whether $f$ satisfies ETE or not is dependent on $q_{o_{2}}$.
We assume that $r_{a_{1}}=\psi^{\prime}$ and this rank-minimizing mechanism is used. First, suppose $q_{o_{2}}=1$. In this case, $\bar{k}\left(\psi\right)=\bar{k}\left(\psi^{\prime}\right)=3$. Then, $r_{a_{1}}=\psi^{\prime}$ is not strategically dominated by $\psi$ and any other preference orders, because
$\left(f\left(\psi^{\prime},\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f\left(\psi^{\prime},\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{2}}=\left(f\left(\psi^{\prime},\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{3}}=\frac{1}{3},$
$\left(f\left(\psi,\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f\left(\psi,\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{2}}=\left(f\left(\psi,\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}\varnothing}=\frac{1}{3}.$
That is, if $p_{a_{2}}=p_{a_{3}}=\psi^{\prime}$, then $a_{1}$ is worse off by revealing $\psi$ instead of $\psi^{\prime}$. Moreover, in this case, $f$ satisfies ETE and thus this is a fair rank-minimizing mechanism.
Second, suppose $q_{o_{2}}\geq 2$. In this case, $\bar{k}\left(\psi\right)=\bar{k}\left(\psi^{\prime}\right)=2$. Then, $a_{1}$ must have either $o_{1}$ or $o_{2}$ if it chooses either $\psi$ or $\psi^{\prime}$ and thus, the unique interest of $a_{1}$ is increasing the probability to obtain $o_{1}$. Thus, in this case, $r_{a_{1}}=\psi^{\prime}$ is strategically dominated by $\psi$. Therefore, even if a fair rank-minimizing mechanism is used, truth-telling may be strategically dominated by some other strategy. Moreover, in this case, $f$ does not satisfy ETE. First, since $q_{o_{1}}+q_{o_{2}}\geq\left|\mathcal{A}\right|$, any agent is assigned to neither $o_{3}$ nor $\varnothing$,∅ , as long as the assignment is not wasteful. Thus, $\psi$ and $\psi^{\prime}$ are essentially the same in this case. However, $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{1}}=0,$ but $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{2}}>0$ or $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{3}}>0$ must be satisfied. Therefore, in this case, $f$ does not satisfy ETE.
4 Refusal option
In this section, we consider the following three-stage game, which is a modification of that in Section 3; that is, the first and second stages in this game are the same as in the previous game. In the third stage, all agents decide whether accept or refuse the object type that is assigned to them. If an agent refuses their assignment, then the agent instead obtains the null object type $\varnothing$ (the outside option). In this game, the null object type has two characteristics: first, there are sufficiently many copies of this type to allocate to all individuals, and second, agents can always obtain it at no cost as long as they abandon their assignment.
This third stage is also considered by Featherstone (2020) but is not explicitly addressed by almost previous studies related to ours. [^3] Doğan and Yenmez (2019) also explicitly consider this stage. However, this stage is implicitly considered by many of them; that is, they consider an ex-post individually rational (or more strictly an ex-post efficient) matching as a requirement and if an assignment is ex-post individually rational, no agent refuses their assignment. See, for example, Bogomolnaia and Moulin (2001) on the property.
Moreover, note that, in this study, all agents are assumed to know the capacities of all object types $\left(q_{o}\right)_{o\in\mathcal{O}}$, when they reveal their preference. On the other hand, Featherstone (2020) considers the situation where any agents do not know the capacity of any object types. Due to this difference, some of our results differ from his.
In the third stage, the obvious optimal (dominant) strategy of an agent $a$ is to accept their assignment denoted by $o$ if $o$ acceptable; that is, if $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$ and refuse it if otherwise ($R\left(o,r_{a}\right)\geq R\left(\varnothing,r_{a}\right)$). Based on this optimal strategy in the third stage, an assignment obtained after the second stage denoted by $x$, and the profile of (true) preferences denoted by $r$, we let $g\left(x,r\right)$ be such that for all $a\in\mathcal{A}$,
$g\left(x,r\right)_{ao}$ $=$ $x_{ao}\text{ for all }R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$
$g\left(x,r\right)_{ao}$ $=$ 0 for all $0\text{ for all }R\left(o,r_{a}\right)>R\left(\varnothing,r_{a}\right)$
$g\left(x,r\right)_{a\varnothing}$ $=$ $\sum\limits_{o:R\left(o,r_{a}\right)\geq R\left(\varnothing,r_{a}\right)}x_{ao}.$
That is, since an agent refuses any of the unacceptable object types when it is assigned, the probability that the null object type is assigned will increase by the probability that any of the unacceptable object types is assigned.
We reconsider strategic problems of a fair rank-minimizing mechanism. For a mechanism $f$, (revealing) $p_{a}^{\prime}$ (for $a$) is said to be strategically dominated by (revealing) $p_{a}$ with the refusal option if $a$ weakly prefers $g\left(f\left(p_{a},p_{-a}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}\right),r\right)$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and strictly prefers $g\left(f\left(p_{a},p_{-a}^{\prime}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}^{\prime}\right),r\right)$ for some $p_{-a}^{\prime}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$.
We show that, under the uniform rank-minimizing mechanism, truth-telling is strategically dominated by another strategy when the agents have the refusal option. We let $\mathcal{D}_{a}\subseteq\mathcal{P}$ be a set of preferences such that $d_{a}\in\mathcal{D}_{a}$ if (i) $R\left(o,r_{a}\right)=R\left(o,d_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$ and (ii) $R\left(\varnothing,d_{a}\right)=\left|\mathcal{O}\right|$. First, the ranks of the truly acceptable object types are unchanged. Second, the null object type $\varnothing$ is the least preferable one under $d_{a}\in\mathcal{D}_{a}$. Thus, any object type that is less preferable than $\varnothing$ in the true preference has changed to be more preferable than $\varnothing$ in $d_{a}$. We call the strategy of agent $a$ of revealing a preference in $\mathcal{D}_{a}$ an outside option demotion strategy (hereafter ODS) of agent $a$. Note that an ODS (revealing $d_{a}\in\mathcal{D}_{a}$) increases the number of acceptable object types revealed by $a$; that is, $R\left(\varnothing,r_{a}\right)\leq R\left(\varnothing,d_{a}\right)=\left|\mathcal{O}\right|$ is satisfied.
Featherstone (2020) defines the full extension of $r_{a}$denoted by $e_{a}$ in this study such that (i) $R\left(o,r_{a}\right)=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, and (ii) $R\left(\varnothing,e_{a}\right)=\left|\mathcal{O}\right|$, and $R\left(o,r_{a}\right)-1=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)>R\left(\varnothing,r_{a}\right)$; that is, $e_{a}\in\mathcal{D}_{a}$ is such that all non-null object types are placed in their true preference orders.
First, we consider the following example to briefly consider the main result.
Example 2
Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$ and $\mathcal{O=}\left\{o_{1},o_{2},\varnothing\right\}$. Suppose $q_{o_{1}}=q_{o_{2}}=1$ and $r_{a_{1}}=(o_{1},\varnothing,o_{2})$. Then, the unique ODS is revealing $d_{a_{1}}=(o_{1},o_{2},\varnothing)$. First, suppose
$p_{a_{2}}=(o_{1},o_{2},\varnothing)\text{ and }p_{a_{3}}=(o_{2},o_{1},\varnothing).$
Then,
$f^{U}\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=y=\left(\begin{array}[]{ccc}y_{a_{1}o_{1}}&y_{a_{1}o_{2}}&y_{a_{1}\varnothing}\\ y_{a_{2}o_{1}}&y_{a_{2}o_{2}}&y_{a_{2}\varnothing}\\ y_{a_{3}o_{1}}&y_{a_{3}o_{2}}&y_{a_{3}\varnothing}\end{array}\right)=\left(\begin{array}[]{ccc}0&0&1\\ 1&0&0\\ 0&1&0\end{array}\right).$
Whenever $\left|\mathcal{A}\right|=\left|\mathcal{O}\right|=3$, we write the matrix representing an assignment in this manner. On the other hand,
$f^{U}\left(d_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\begin{array}[]{ccc}1/2&0&1/2\\ 1/2&0&1/2\\ 0&1&0\end{array}\right)$
Thus, revealing $d_{a_{1}}$ instead of $r_{a_{1}}$ may increase the probability that $a_{1}$ is assigned to their best object $o_{1}$. On the other hand, suppose
$p_{a_{2}}^{\prime}=(o_{1},\varnothing,o_{2})\text{ and }p_{a_{3}}^{\prime}=(o_{1},\varnothing,o_{2}).$
Then,
$f^{U}\left(r_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&0&2/3\\ 1/3&0&2/3\end{array}\right),$
$f^{U}\left(d_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)=\left(\begin{array}[]{ccc}1/3&2/3&0\\ 1/3&0&2/3\\ 1/3&0&2/3\end{array}\right).$
Thus, in this case, $a_{1}$ seems to be worse off by revealing $d_{a_{1}}$ instead of $r_{a_{1}}$, because $a_{1}$ prefers $\varnothing$ to $o_{2}$. However, in the model introduced in this section, $a_{1}$ can refuse $o_{2}$ and in that case, $\varnothing$ is assigned. Therefore,
$g\left(f^{U}\left(d_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right),r\right)=g\left(f^{U}\left(r_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right),r\right)\text{;}$
that is, $a_{1}$ is not worse off by choosing $d_{a_{1}}$ instead of $r_{a_{1}}$.
Next, let $\hat{p}_{a_{2}}=\hat{p}_{a_{3}}=(o_{1},o_{2},\varnothing)$. Then,
$f^{U}\left(r_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right)=\left(\begin{array}[]{ccc}0&0&1\\ 1/2&1/2&0\\ 1/2&1/2&0\end{array}\right),$
$f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right)=\left(\begin{array}[]{ccc}1/3&1/3&1/3\\ 1/3&1/3&1/3\\ 1/3&1/3&1/3\end{array}\right).$
In the latter case, $a_{1}$ may refuse the assignment; that is,
$g\left(f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right),r\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&1/3&1/3\\ 1/3&1/3&1/3\end{array}\right),$
which is wasteful, because
$\sum\limits_{a\in\left\{a_{1},a_{2},a_{3}\right\}}g\left(f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right),r\right)_{ao_{2}}$ $=$ $\frac{2}{3}<1,$
and $\text{and }g\left(f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right),r\right)_{a^{\prime}\varnothing}$ $=$ $\frac{1}{3}>0,$
for $a^{\prime}=a_{2},a_{3}$. Thus, it is ordinally dominated by
$\tilde{x}=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&1/2&1/6\\ 1/3&1/2&1/6\end{array}\right).$
In the case where $a_{1}$ is assigned to $o_{2}$, either $a_{2}$ or $a_{3}$ is assigned to $\varnothing$ but $a_{1}$ will refuse $o_{2}$. Thus, in that case, it is more efficient that the agent who is assigned to $\varnothing$ is assigned to $o_{2}$ and $a_{1}$ is assigned to $\varnothing$. That is, revealing $d_{a_{1}}$ and refusing some unacceptable assignment may make $a_{1}$ better off but result in a wasteful assignment, because some agents may desire the object type refused by $a_{1}$.
Note that $\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right)$ can be a Nash equilibrium when $r_{a_{2}}=\hat{p}_{a_{2}}$ and $r_{a_{3}}=\hat{p}_{a_{3}}$. Trivially, $a_{1}$ has no incentive to change $d_{a_{1}}$. On the other hand, suppose that both $a_{2}$ and $a_{3}$ prefer the assignment such that each object $o\in\left\{o_{1},o_{2},\varnothing\right\}$ assigns with the probability $1/3$ to the assignment such that $o_{2}$ assigns with probability $1$. This ensures that neither $a_{2}$ nor $a_{3}$ has an incentive to change $\hat{p}_{a_{2}}$ and $\hat{p}_{a_{3}}$ to $(o_{2},o_{1},\varnothing)$ or $(o_{2},\varnothing,o_{1})$. Then, in this case, $\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right)$ is a Nash equilibrium. Thus, a wasteful assignment can occur as a Nash equilibrium of agents’ strategies when the assignment is determined by the uniform rank-minimizing mechanism.
Now, we formally introduce our main results.
Theorem 1
Suppose that the uniform rank-minimizing mechanism used. Then, for all agents, truth-telling is weakly strategically dominated by any ODS with the refusal option.
By this result, an agent is never hurt by using an ODS instead of truth-telling. Next, we show that with some conditions, there exists some ODS such that an agent may be better off by using it instead of truth-telling.
Theorem 2
Suppose that there is $a\in\mathcal{A}$ whose true preference $r_{a}$ satisfies
$R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)<R\left(o^{\prime},r_{a}\right)\text{ and }\sum\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}<\left|\mathcal{A}\right|\text{,}$ (3)
[…]
This result implies that if there is $a\in\mathcal{A}$ satisfying (3) for some $o,o^{\prime}\in\mathcal{O}$, then agent $a$ has a strong incentive to adopt an ODS. This is satisfied only when some agent has at least one unacceptable object. The condition given in (3) is that there is a truly unacceptable type $o^{\prime}$ such that the number of copies of $o^{\prime}$, plus the sum of the copies of the truly acceptable types, is sufficiently small. We discuss the cases where (3) is satisfied for some $a,o,o^{\prime}$ in the subsequent section. Note that there is $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $R\left(\varnothing,p_{a^{\prime}}\right)=\left|\mathcal{O}\right|$ for all $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$ and $a$ strictly prefers $g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$. Thus, for all $a^{\prime}$, $p_{a^{\prime}}$ may also be an ODS of $a^{\prime}$.
Theorems 1 and 2 do not say that an ODS is a dominant strategy. However, when if there is $a\in\mathcal{A}$ whose true preference satisfies (3) for some $o,o^{\prime}\in\mathcal{O}$, since truth-telling is strategically dominated by another strategy, $a$ has a strong incentive to manipulate its preference ranking, while one of the ODSs may not necessarily be chosen.
As stated earlier, there may exist multiple ODSs. In that case, an ODS may be strategically dominated by another one. We have the following result.
Proposition 3
Suppose that there is $a\in\mathcal{A}$ whose true preference $r_{a}$ satisfies
$R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)<R\left(o^{\prime},r_{a}\right)\text{ and }\sum\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}\geq\left|\mathcal{A}\right|,$
for some $o,o^{\prime}\in\mathcal{O}$. Let $d_{a}\in\mathcal{D}_{a}$ be such that $R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},d_{a}\right)$. Then, for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and all $\bar{o}\in\mathcal{O}$,
$g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)_{a\bar{o}}=g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)_{a\bar{o}}.$
In this case, truth-telling is not strategically dominated by an ODS of revealing such $d_{a}$. Note that revealing the full extension may be such an ODS, when $o^{\prime}$ defined in Proposition 3 satisfies $R\left(\varnothing,r_{a}\right)+1=R\left(o^{\prime},r_{a}\right)$.
Moreover, by Theorem 1 and Proposition 3, we show that for an agent, revealing the full extension may be strategically dominated by another ODS. First, we provide an example to briefly see this fact.
Example 3
Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$ and $\mathcal{O=}\left\{o_{1},o_{2},o_{3},\varnothing\right\}$. Suppose $q_{o_{1}}=q_{o_{3}}=1,$ $q_{o_{2}}=2$ and $r_{a_{1}}=(o_{1},\varnothing,o_{2},o_{3})$. Then, there are two ODSs. One is $e_{a_{1}}=(o_{1},o_{2},o_{3},\varnothing)$, which is the full extension of $a_{1}$, and the other is $d_{a_{1}}=(o_{1},o_{3},o_{2},\varnothing)$. First, suppose
$\bar{p}_{a_{2}}=(o_{1},o_{3},o_{2},\varnothing)\text{ and }\bar{p}_{a_{3}}=(o_{3},o_{1},o_{2},\varnothing).$
Then,
$f^{U}\left(r_{a_{1}},\bar{p}_{a_{2}},\bar{p}_{a_{3}}\right)=y=\left(\begin{array}[]{cccc}y_{a_{1}o_{1}}&y_{a_{1}o_{2}}&y_{a_{1}o_{3}}&y_{a_{1}\varnothing}\\ y_{a_{2}o_{1}}&y_{a_{2}o_{2}}&y_{a_{2}o_{3}}&y_{a_{2}\varnothing}\\ y_{a_{3}o_{1}}&y_{a_{3}o_{2}}&y_{a_{3}o_{3}}&y_{a_{3}\varnothing}\end{array}\right)=\left(\begin{array}[]{cccc}0&0&0&1\\ 1&0&0&0\\ 0&0&1&0\end{array}\right),$
In this example, we write the assignments in this manner. Moreover,
$g\left(f^{U}\left(d_{a_{1}},\bar{p}_{a_{2}},\bar{p}_{a_{3}}\right),r\right)=\left(\begin{array}[]{cccc}1/2&0&0&1/2\\ 1/2&1/2&0&0\\ 0&0&1&0\end{array}\right)$
Therefore, $a_{1}$ (strictly) prefers $g\left(f^{U}\left(d_{a_{1}},\bar{p}_{a_{2}},\bar{p}_{a_{3}}\right),r\right)$ to $g\left(f^{U}\left(r_{a_{1}},\bar{p}_{a_{2}},\bar{p}_{a_{3}}\right),r\right)$. On the other hand, there is no pair $\left(p_{a_{2}},p_{a_{3}}\right)$ of preference orders of $a_{2}$ and $a_{3}$ such that $a_{1}$ (strictly) prefers $g\left(f^{U}\left(e_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)$ to $g\left(f^{U}\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)$. This is because, by Proposition 3 and $q_{o_{1}}+q_{o_{2}}=3=\left|\mathcal{A}\right|$, for any $\left(p_{a_{2}},p_{a_{3}}\right)\in\mathcal{P}^{2}$ and any $\bar{o}\in\mathcal{O}$,
$\left(g\left(f^{U}\left(e_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)\right)_{a_{1}\bar{o}}=\left(g\left(f^{U}\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)\right)_{a_{1}\bar{o}}\text{.}$
We explain why $d_{a_{1}}$ strategically dominates $e_{a_{1}}$. In this example, $q_{o_{1}}+q_{o_{3}}=2<3=q_{o_{1}}+q_{o_{2}}$. That is, the copies of $o_{1}$ and $o_{2}$ are sufficient to allocate all agents, but those of $o_{1}$ and $o_{3}$ are insufficient. Thus, revealing $o_{2}$ as a second best type does not increase the probability that $a_{1}$ is assigned to $o_{1}$. On the other hand, revealing $o_{3}$ as a second best type increases it, when $o_{3}$ is the first best type of another agent. Thus, revealing the full extension may be strategically dominated by another ODS when an unacceptable type (in this case $o_{3}$) is less preferable than the other (in this case $o_{2}$) but the number of copies of the former is less than that of the latter.
Generally, we have the following result.
Lemma 4
Suppose that there is $a\in\mathcal{A}$ whose true preference satisfies
$R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},r_{a}\right)-1<R\left(o^{\prime\prime},r_{a}\right),\text{ }$
$\sum\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}\geq\left|\mathcal{A}\right|,\text{ and}$
$\sum\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime\prime}}<\left|\mathcal{A}\right|,$
[…]
$R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},e_{a}\right)=R\left(o^{\prime\prime},d_{a}\right),$
with the refusal option.
Note that Lemma 4 does not contradict a result of Featherstone (2020, Proposition 13). In his model, he assumes that the agents have no information on the capacities of object types. Revealing the full extension can be strategically dominated by another ODS only when the capacities of the unacceptable types are different and agents know them. Thus, the strategic domination in Lemma 4 does not occur in the low-information situation discussed by Featherstone (2020).
Next, we show that if agents adopt an ODS, then the result may be wasteful and thus inefficient.
Proposition 4
Suppose that there is $a\in\mathcal{A}$ whose true preference satisfies (3) for some $o,o^{\prime}\in\mathcal{O}$. Then, there are some ODS of revealing $d_{a}$ and some $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $R\left(\varnothing,p_{a^{\prime}}\right)=\left|\mathcal{O}\right|$ for all $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$ and $g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)$ is wasteful.
By Theorem 2, for an agent $a$ with $r_{a}$ that satisfy (3) for some $o,o^{\prime}\in\mathcal{O}$, an ODS dominates truth-telling. Moreover, by Proposition 1, in the same case, the strategy may make the assignment wasteful and thus inefficient. In an ODS, an agent reveals actually unacceptable types as acceptable types. Then, the agent may be assigned to an actually unacceptable type and refuse it. Since the type may be desirable for another agent, in that case, the assignment is wasteful.
By Theorems 1 and 2 and Proposition 3, if the uniform rank-minimizing mechanism is used, the agents may have a strong incentive to manipulate their preferences, and these manipulations may result in an inefficient assignment. Thus, next, we consider another fair rank-minimizing mechanism such that, contrary to $f^{U}$, if it is used as the mechanism in the second stage, then for any agent, truth-telling is not strategically dominated by any other strategy.
We let $f^{M}$ be the modified uniform rank-minimizing mechanism such that if $p$ satisfies
for agent $a,$ $R\left(o,p_{a}\right)=1$ and $R\left(\varnothing,p_{a}\right)\geq 3,$
there are $q_{o}$ agents represented by $a^{\prime}$ revealing $p_{a^{\prime}}$, such that $R\left(\varnothing,p_{a^{\prime}}\right)=l\geq 2,$ $p_{a}^{l^{\prime}}=p_{a^{\prime}}^{l}$ for all $l^{\prime}=1,\cdots,l-1,$ $R\left(\varnothing,p_{a}\right)>l,$ and $\bar{k}(p_{a^{\prime}})=l$; that is,
$\sum\limits_{\bar{o}:R\left(\bar{o},p_{a^{\prime}}\right)<R\left(\varnothing,p_{a^{\prime}}\right)}q_{\bar{o}}<\left|\mathcal{A}\right|,$
if $\left|\mathcal{A}\right|>q_{o}+1$, then the first best type of the other agents ($\left|\mathcal{A}\right|-\left(q_{o}+1\right)$ agents) represented by $a^{\prime\prime}$ is $\varnothing$, then
$\left(f^{M}\left(p\right)\right)_{ao}=0,\text{ }\left(f^{M}\left(p\right)\right)_{a^{\prime}o}=1,\text{ }\left(f^{M}\left(p\right)\right)_{ap_{a}^{2}}=1\text{ and }\left(f^{M}\left(p\right)\right)_{a^{\prime\prime}\varnothing}=1,$
and if $p$ does not satisfy all of them, then
$f^{M}\left(p\right)=\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}=f^{U}\left(p\right).$
To understand this mechanism, we reconsider Example 1. First, suppose $q_{o_{1}}=q_{o_{2}}=1$. If $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\psi^{\prime},\psi,\psi\right)$, $\left(\psi,\psi^{\prime},\psi\right)$ or $\left(\psi,\psi,\psi^{\prime}\right)$, then this satisfies (i), (ii) and (iii). In these cases, a student who chooses $\psi^{\prime}$ is corresponding to $a,$ another student is corresponding to $a^{\prime}$, and $o_{1}$ is corresponding to $o$. The rank-minimizing mechanism $f$ introduced in Example 1 is the modified uniform rank-minimizing mechanism, because a student who chooses $\psi^{\prime}$ obtains $o_{1}$ with probability 00 and another student obtains $o_{1}$ with probability $1$
On the other hand, suppose $q_{o_{1}}=1$ and $q_{o_{2}}\geq 2$. Even if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\psi^{\prime},\psi,\psi\right)$, $\left(\psi,\psi^{\prime},\psi\right)$ or $\left(\psi,\psi,\psi^{\prime}\right)$, this does not satisfy (ii), because in this case, $p_{a^{\prime}}=\psi$ and $l=3.$ Then,
$\sum\limits_{\bar{o}:R\left(\bar{o},p_{a^{\prime}}\right)<R\left(\varnothing,p_{a^{\prime}}\right)}q_{\bar{o}}=q_{o_{1}}+q_{o_{2}}=3=\left|\mathcal{A}\right|.$
Therefore, (ii) is not satisfied with $p_{a^{\prime}}=\psi$.
We have the following result.
Proposition 5
The modified uniform rank-minimizing mechanism $f^{M}$ satisfies ETE.
The strategic issues, mentioned in Theorem 2, do not arise under this mechanism $f^{M}$.
First, we explain why the strategic issues mentioned in Example 2 does not arise in the modified uniform rank-minimizing mechanism. If $r_{a}=p_{a^{\prime}}$ where $p_{a^{\prime}}$ is defined in (ii), then $\left(d_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii) where $d_{a}\in\mathcal{D}_{a}$. This implies that for some $p_{-a}$, if agent $a$ adopts an ODS instead of truth-telling and $f^{M}$ is used, then $a$ is worse off.
To see this fact, we revisit Example 2. In this example, $\left(d_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)$ satisfies (i), (ii) and (iii), but $\left(r_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)$ does not. Thus,
$f^{M}\left(r_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)$ $=$ $f^{U}\left(r_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&0&2/3\\ 1/3&0&2/3\end{array}\right),$
$f^{M}\left(d_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)$ $=$ $\left(\begin{array}[]{ccc}0&1&0\\ 1/2&0&1/2\\ 1/2&0&1/2\end{array}\right),$
$g\left(f^{M}\left(d_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)\right)$ $=$ $\left(\begin{array}[]{ccc}0&0&1\\ 1/2&0&1/2\\ 1/2&0&1/2\end{array}\right).$
This implies that $a_{1}$ may be worse off by revealing $d_{a_{1}}$ instead of $r_{a_{1}}$, even if $a_{1}$ refuses any unacceptable objects. Therefore, $r_{a_{1}}$ is not strategically dominated by $d_{a_{1}}$, if $f^{M}$ is used.
Formally, we have the following fact.
Proposition 6
Suppose that the modified uniform rank-minimizing mechanism is used. For any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ that strategically dominates $r_{a}$ with (or without) the refusal option.
By adopting $f^{M}$, any ODS can fail, because increasing the number of acceptable objects may lead the agent to be unassigned acceptable objects in their true preference. Thus, in some profiles of other agents’ (revealed) preferences, truth-telling is better than adopting any ODS. However, $f^{M}$ may lead to other strategic problems. That is, if $f^{M}$ is adopted instead of $f^{U}$, an agent may have an incentive to reduce the number of acceptable object types.
Example 4
Let $R\left(\varnothing,r_{a}\right)=k$. Suppose $k\geq 3$; that is, $\varnothing$ is neither the best type nor the second best type for $r_{a}$. Thus, we let $r_{a}^{1}=o$ and $r_{a}^{2}=o^{\prime}$. For $p_{-a}$, there are $q_{o}$ agents represented by $a^{\prime}$ whose revealed preference order is such that $p_{a^{\prime}}^{k^{\prime}}=r_{a^{\prime}}^{k^{\prime}}$ for all $k^{\prime}=1,\cdots,k-2$ and $p_{a^{\prime}}^{k-1}=\varnothing$. We assume
$q_{r_{a}^{1}}+\cdots+q_{r_{a}^{k-2}}<\left|\mathcal{A}\right|\text{.}$
Moreover, for $p_{-a}$ and any other agent represented by $a^{\prime\prime}$, $R\left(\varnothing,r_{a^{\prime\prime}}\right)=1$. In this case, since $\left(r_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii),
$\left(f^{M}\left(r_{a},p_{-a}\right)\right)_{ao}$ $=$ 0, ($0,\text{ }\left(f^{M}\left(r_{a},p_{-a}\right)\right)_{ao^{\prime}}=1,$
$\left(f^{U}\left(r_{a},p_{-a}\right)\right)_{ao}$ $=$ $\frac{q_{o}}{q_{o}+1},$
$\left(f^{U}\left(r_{a},p_{-a}\right)\right)_{ao^{\prime}}$ $=$ $\frac{1}{q_{o}+1}.$
On the other hand, we let $p_{a}=p_{a^{\prime}}$. Since $\left(p_{a},p_{-a}\right)$ does not satisfy all of (i), (ii) and (iii),
$\left(f^{M}\left(p_{a},p_{-a}\right)\right)_{ao}$ $=$ $\left(f^{U}\left(p_{a},p_{-a}\right)\right)_{ao}=\frac{q_{o}}{q_{o}+1},$
$\left(f^{M}\left(p_{a},p_{-a}\right)\right)_{ao^{\prime}}$ $=$ $\left(f^{U}\left(p_{a},p_{-a}\right)\right)_{ao^{\prime}}=\frac{1}{q_{o}+1}.$
Thus, in this case, if $f^{M}$ is used, then $a$ has an incentive to choose $p_{a}$. In other words, in this case, $a$ has an incentive to reveal by reducing its truly acceptable types. On the other hand, if $f^{U}$ is used, then $a$ has no such an incentive.
Of course, this does not mean that truth-telling is strategically dominated revealing some reduction of $r_{a}$. However, agents may have such incentives even if they do not have the refusal option.
5 Concluding Remarks
We show that if an assignment is determined by the uniform rank-minimizing mechanism, which is a fair one, and agents can refuse the assignment and obtain the outside option instead, then the true preference is strategically dominated by an ODS. The adoption of such strategies by agents may result in inefficient assignments; that is, an agent who adopts an ODS refuses their assigned type, although some other agents desire that type.
We consider a public school choice system as an example. Our result implies that a student may have an incentive to reveal truly unacceptable schools as acceptable schools, and they may be assigned to one of them and then refuse the admission even if some other students desire the school. In such a case, the school may re-enroll students, but even in that case, it would cause market disruption.
Our result implies that if all agents prefer many other types over the outside option, then the strategic problem considered in this study is not so severe. For example, in the case of matching systems for schools that are part of compulsory education, the outside option does not need to be considered. Therefore, there is no strategic problem of the kind discussed in this study. On the other hand, the strategic problem may be severe when there are agents who relatively highly value the outside option.
Moreover, if agents who refuse their assignment can be sufficiently penalized, the problem would be resolved. However, considering that some agents may refuse their assignment due to unavoidable circumstances (without any strategic intention), the penalties should not be too severe.
Therefore, if resolving the problem in that way is difficult, it is necessary to consider alternative mechanisms. First, as shown in this study, if we modify the uniform rank-minimizing mechanism, then truth-telling is not strategically dominated by any other strategies, including the outside option demotion ones. However, in this modification, it is necessary to ensure that increasing the number of acceptable types can result in a disadvantage. Thus, this modification may lead to the opposite problem, where agents strategically decrease the number of acceptable types. Second, we consider efficient and fair mechanisms other than the fair rank-minimizing mechanisms, such as the (general) probabilistic serial mechanism introduced by Bogomolnaia and Moulin (2001) and Budish et al. (2013), because it is fair and ordinally efficient, although as stated earlier, this efficiency property is weaker than the rank efficiency.
Appendix
Proof of Lemma 1
First, by Remark 1, any $x^{\ast}\in\mathcal{X\setminus Y}$ can be represented by a convex combination of deterministic assignments $y^{1},\cdots,y^{I}$ with $I\geq 2$. Suppose not; that is, there is $y^{i}$ such that $RV\left(y^{i},p\right)>RV\left(x^{\ast},p\right)$.
$RV\left(x^{\ast},p\right)$ $=$ $\sum\limits_{a\in\mathcal{A}}\sum\limits_{o\in\mathcal{O}}R\left(o,p_{a}\right)\times x_{ao}^{\ast}$ (7)
$=$ $\sum\limits_{a\in\mathcal{A}}\sum\limits_{o\in\mathcal{O}}\left(R\left(o,p_{a}\right)\times\sum\limits_{i=1}^{I}\sigma^{i}y_{ao}^{i}\right)$
$=$ $\sum\limits_{i=1}^{I}\sigma^{i}\sum\limits_{a\in\mathcal{A}}\sum\limits_{o\in\mathcal{O}}R\left(o,p_{a}\right)\times y_{ao}^{i}$
$=$ $\sum\limits_{i=1}^{I}\sigma^{i}RV\left(y^{i},p\right).$
Thus, $RV\left(y^{i},p\right)>RV\left(x^{\ast},p\right)$ implies that there is $y^{j}$ such that $RV\left(y^{j},p\right)<RV\left(x^{\ast},p\right)$, which contradicts that $x^{\ast}$ is a rank efficient assignment.
Proof of Lemma 2
Fix $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ and $a\in\mathcal{A}$. Suppose not; that is, for an assignment $x\in\mathcal{X}$ that is not wasteful for $p$, there is $o\in\mathcal{O}$ satisfying $R(o,p_{a})>\bar{k}\left(p_{a}\right)$ and $x_{ao}>0$. First, since $\sum_{\bar{o}\in\mathcal{O}}x_{a\bar{o}}=1$,
$\sum\limits_{\bar{o}\in\mathcal{O}}\sum\limits_{\bar{a}\in\mathcal{A}}x_{\bar{a}\bar{o}}=\left|\mathcal{A}\right|\text{.}$
Moreover, since $x_{ao}>0$,
$\sum\limits_{\bar{o}:R(\bar{o},p_{a})\leq\bar{k}\left(p_{a}\right)}\sum\limits_{\bar{a}\in\mathcal{A}}x_{\bar{a}o^{\prime}}$ $<$ $\left|\mathcal{A}\right|\leq\sum\limits_{\bar{o}:R(\bar{o},p_{a})\leq\bar{k}\left(p_{a}\right)}q_{o^{\prime}}\text{, and}$
$\sum\limits_{\bar{o}:R(\bar{o},p_{a})\leq\bar{k}\left(p_{a}\right)}x_{a\bar{o}}$ $<$ $1$
Since $\sum_{\bar{a}\in\mathcal{A}}x_{\bar{a}o^{\prime\prime}}\leq q_{o^{\prime\prime}}$ for all $o^{\prime\prime}\in\mathcal{O}$, there is some $o^{\prime}\in\mathcal{O}$ such that $R(o^{\prime},p_{a})\leq\bar{k}\left(p_{a}\right)$ and $\sum_{\bar{a}\in\mathcal{A}}x_{\bar{a}o^{\prime}}<q_{o^{\prime}}$. Then, $\sum_{\bar{a}\in\mathcal{A}}x_{\bar{a}o^{\prime}}<q_{o^{\prime}},$ $x_{ao}>0,$ and $R(o^{\prime},p_{a})\leq\bar{k}\left(p_{a}\right)<R(o,p_{a})$ contradict that $x\in\mathcal{X}$ is not wasteful for $p$.
Proof of Proposition 1
Let $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfy $p_{a}^{k^{\prime}}=p_{a^{\prime}}^{k^{\prime}}$ for all $1,\cdots,\bar{k}\left(p_{a}\right)$. Suppose not; that is, $\left(f^{U}\left(p\right)\right)_{ao}>\left(f^{U}\left(p\right)\right)_{a^{\prime}o}$ for some $o\in\mathcal{O}$. Since $\left(f^{U}\left(p\right)\right)_{ao}>0$ and $f^{U}\left(p\right)$ is not wasteful, $R\left(o,p_{a}\right)\leq\bar{k}\left(p_{a}\right)$, by Lemma 2. Therefore, $R\left(o,p_{a}\right)=R\left(o,p_{a^{\prime}}\right)$.
Now, let
$Y=\left\{y\in\mathcal{Y}^{\ast}\left(p\right)\left|\text{ }y_{ao}=1\right.\right\},Y^{\prime}=\left\{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)\left|\text{ }y_{a^{\prime}o}^{\ast}=1\right.\right\}.$
Since $\left(f^{U}\left(p\right)\right)_{ao}>0$, $Y\neq\emptyset$ and thus let $Y=\left\{y^{1},\cdots,y^{\left|Y\right|}\right\}$. Then, for each $i=1,\cdots,\left|Y\right|$, there is $y^{i\prime}$ be such that $y_{ao}^{i\prime}=0,$ $y_{a^{\prime}o}^{i\prime}=1,$ and $y_{a^{\prime\prime}o^{\prime}}^{i\prime}=y_{a^{\prime\prime}o^{\prime}}^{i}$ for all $o^{\prime}\in\mathcal{O}\setminus\{o\}$ and all $a^{\prime\prime}\in\mathcal{A}\setminus\{a,a^{\prime}\}$. Since $R\left(o,p_{a}\right)=R\left(o,p_{a^{\prime}}\right)$, $RV\left(y^{i},p\right)=RV\left(y^{i\prime},p\right)$ and therefore $y^{i\prime}\in Y^{\prime}$. Moreover, since $y^{1\prime},$ $\cdots,y^{\left|Y\right|\prime}$ are distinct, $\left|Y\right|\leq\left|Y^{\prime}\right|$. However,
$\left(f^{U}\left(p\right)\right)_{ao}=\frac{\left|Y\right|}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}\leq\frac{\left|Y^{\prime}\right|}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}=\left(f^{U}\left(p\right)\right)_{a^{\prime}o}$
contradicts $\left(f^{U}\left(p\right)\right)_{ao}>\left(f^{U}\left(p\right)\right)_{a^{\prime}o}$.
Proof of Lemma 3
Suppose $k\leq\bar{k}\left(p_{a}\right)$. Then, since $q_{\varnothing}\geq\left|\mathcal{A}\right|$, $r_{a}^{k^{\prime}}=p_{a}^{k^{\prime}}\neq\varnothing$ for all $k^{\prime}=1,\cdots,k-1$.
When $k=1$, we let $R(\varnothing,p_{a^{\prime}})=1$ for all $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$. Then,
$\left(f\left(r_{a},p_{-a}\right)\right)_{ar_{a}^{1}}=\left(f\left(p_{a},p_{-a}\right)\right)_{ap_{a}^{1}}=1,$
where $R(r_{a}^{1},r_{a})=1<R(p_{a}^{1},r_{a})$. Thus, $a$ with $r_{a}$ strictly prefers $f\left(r_{a},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$.
Next, we assume $k=2$. Then, $p_{a}^{1}=r_{a}^{1}$ and $p_{a}^{2}\neq r_{a}^{2}$. We assume that under $\left(r_{a},p_{-a}\right)$, the agents are divided into two groups. For the first group, there are $q_{r_{a}^{1}}+1$ agents whose preference is equivalent to $r_{a}$ under $\left(r_{a},p_{-a}\right)$; that is, if $a^{\prime}$ belongs to this group, then $r_{a}=p_{a^{\prime}}$. Note that $a$ itself belongs to this group. For the second group, the best object type is $\varnothing$ under $\left(r_{a},p_{-a}\right)$. Then, since all agents in the second group always assign $\varnothing$ with probability $1$ and $f$ satisfies ETE,
$\left(f\left(r_{a},p_{-a}\right)\right)_{ar_{a}^{1}}$ $=$ $\frac{q_{r_{a}^{1}}}{q_{r_{a}^{1}}+1},$
$\left(f\left(r_{a},p_{-a}\right)\right)_{ar_{a}^{2}}$ $=$ $\frac{1}{q_{r_{a}^{1}}+1}.$
On the other hand,
$\left(f\left(p_{a},p_{-a}\right)\right)_{ap_{a}^{1}}$ $=$ $\frac{q_{r_{a}^{1}}}{q_{r_{a}^{1}}+1},$
$\left(f\left(p_{a},p_{-a}\right)\right)_{ap_{a}^{2}}$ $=$ $\frac{1}{q_{r_{a}^{1}}+1}.$
Since $p_{a}^{1}=r_{a}^{1}$ and $p_{a}^{2}\neq r_{a}^{2}$, $a$ strictly prefers $f\left(r_{a},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$.
We consider an arbitrary integer $k\geq 3$. Then, $r_{a}^{k^{\prime}}=p_{a}^{k^{\prime}}$ for all $k^{\prime}=1,\cdots,k-1$. We assume that under $\left(r_{a},p_{-a}\right)$ the agents are divided into $k$ groups. For the first group, there are $q_{r_{a}^{1}}+1$ agents whose preference is equivalent to $r_{a}$ under $\left(r_{a},p_{-a}\right)$; that is, if $a^{\prime}$ is in this group, then $r_{a}=p_{a^{\prime}}$. Note that $a$ belongs to this group. Next, we consider $j$th group for each $j=2,3,\cdots,k-1$. In $j$th group, there are just $q_{r_{a}^{j}}$ agents whose best object type is $r_{a}^{j}$ and the second best object is $r_{a}^{j+1},$ the third best $r_{a}^{j+2},$cdots and the $k$th best $r_{a}^{k}$ under $\left(r_{a},p_{-a}\right)$. If $a_{j}$ belongs $j$th group for $j=1,2,\cdots,k-1$, the preference orders of them are as follows:
$p_{a_{1}}$ $=$ $\left(r_{a}^{1},r_{a}^{2},\ldots,r_{a}^{k-2},r_{a}^{k-1},r_{a}^{k},\ldots\right),$ (8)
$p_{a_{2}}$ $=$ $\left(r_{a}^{2},r_{a}^{3},\ldots,r_{a}^{k-1},r_{a}^{1},r_{a}^{k},,\ldots\right),$
$\vdots$
$p_{a_{j}}$ $=$ $\left(r_{a}^{j},r_{a}^{j+1},\ldots,r_{a}^{j-2},r_{a}^{j-1},r_{a}^{k},\ldots\right),$
$\vdots$
$p_{a_{k-1}}$ $=$ $\left(r_{a}^{k-1},r_{a}^{1},\ldots,r_{a}^{k-3},r_{a}^{k-2},r_{a}^{k},\ldots\right).$
Since $k\leq\bar{k}\left(p_{a}\right)$,
$q_{r_{a}^{1}}+q_{r_{a}^{2}}+\cdots+q_{r_{a}^{k-1}}=\sum\limits_{o:R(o,r_{a})<k}q_{o}<\left|\mathcal{A}\right|.$
and thus $\left(r_{a},p_{-a}\right)$ is well-defined so far. Finally, for all the remaining agents (belonging to the $k$th group), the best object type is $\varnothing$ under $\left(r_{a},p_{-a}\right)$.
In this case, at least one agent cannot be assigned to their best object. Thus, in a rank-minimizing assignment, all agents in $j$th group for all $j=2,3,\cdots,k-1$ is assigned to their best object type $r_{a}^{j}$, and $q_{r_{a}^{1}}$ agents in the first group is also assigned to their best object type $r_{a}^{1},$ and moreover, just one agent in the first group is also assigned to $r_{a}^{k}$. Moreover, an agent who belongs to the $k$th group is assigned to their best object type $\varnothing$.
Since all agents excepting those belonging to the first group are assigned to their best alternative with probability $1$ and $f^{U}$ satisfies ETE,
$\left(f\left(r_{a},p_{-a}\right)\right)_{ar_{a}^{1}}$ $=$ $\frac{q_{r_{a}^{1}}}{q_{r_{a}^{1}}+1},$
$\left(f\left(r_{a},p_{-a}\right)\right)_{ar_{a}^{k}}$ $=$ $\frac{1}{q_{r_{a}^{1}}+1}.$
On the other hand,
$\left(f\left(p_{a},p_{-a}\right)\right)_{ap_{a}^{1}}$ $=$ $\frac{q_{r_{a}^{1}}}{q_{r_{a}^{1}}+1},$
$\left(f\left(p_{a},p_{-a}\right)\right)_{ap_{a}^{k}}$ $=$ $\frac{1}{q_{r_{a}^{1}}+1}.$
Since $p_{a}^{k^{\prime}}=r_{a}^{k^{\prime}}$ for all $k^{\prime}=1,\cdots,k-1$ and $p_{a}^{k}\neq r_{a}^{k}$, $a$ strictly prefers $f^{U}\left(r_{a},p_{-a}\right)$ to $f^{U}\left(p_{a},p_{-a}\right)$. Thus, we have the first result.
Proof of Proposition 2
Let $f$ be a fair mechanism, $p_{a}\neq r_{a}$ and $k$ be the smallest integer such that $p_{ak}\neq r_{ak}$. First, suppose $k\leq\bar{k}\left(p_{a}\right)$. By Lemma 2, there is $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a$ strictly prefers $f\left(r_{a},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$. Therefore, in this case, $r_{a}$ is not strategically dominate by $p_{a}$.
Second, suppose $k>\bar{k}\left(p_{a}\right)$. In this case, since $f$ satisfies ETE, $f\left(r_{a},p_{-a}\right)=f\left(p_{a},p_{-a}\right)$. Therefore, $p_{a}$ does not strategically dominate $r_{a}$.
Proof of Theorem 1
Fix any $\left(r_{a},p_{-a}\right)\in\mathcal{P}^{\left|\mathcal{A}\right|}$ and any $d_{a}\in\mathcal{D}_{a}$. We show that $a$ weakly prefers $g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$. First, suppose that there is no $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}=1$ for some $o\in\mathcal{O}$ with $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$. Then, $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)_{a\varnothing}=1.$ Thus, for any assignment $x$, $a$ weakly prefers $g\left(x,r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$.
Second, suppose that there is some $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}=1$ with $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$. In this case, we have the following two results.
Claim 1
If $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}=1$ where $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, then $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
Proof. By the construction,
$RV\left(y,\left(r_{a},p_{-a}\right)\right)=RV\left(y,\left(d_{a},p_{-a}\right)\right).$
Suppose $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}=1$ where $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$. Toward a contradiction, suppose $y\notin\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
We first consider the case where there is $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$ that satisfies $y_{ao^{\prime}}^{\prime}=1$ for $o^{\prime}$ such that $R\left(o^{\prime},r_{a}\right)<R\left(\varnothing,r_{a}\right)$. By the construction of $d_{a}$,
$RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)$ $=$ $RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)<$
$RV\left(y,\left(d_{a},p_{-a}\right)\right)$ $=$ $RV\left(y,\left(r_{a},p_{-a}\right)\right),$
which contradicts $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$.
We second consider the case where there is no such a deterministic assignment. Since $\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\neq\emptyset$ (by Corollary 1), there is $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$ such that $y_{ao^{\prime}}^{\prime}=1$ for $o$ such that $R\left(o,r_{a}\right)\geq R\left(\varnothing,r_{a}\right)$. In this case,
$RV\left(y,\left(r_{a},p_{-a}\right)\right)=RV\left(y,\left(d_{a},p_{-a}\right)\right)>RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right).$
Now, we consider $y^{\prime\prime}$ such that $y_{a\varnothing}^{\prime\prime}=1,$ $y_{a\bar{o}}^{\prime\prime}=0$ for all $\bar{o}\in\mathcal{O}\setminus\left\{\varnothing\right\}$ and $y_{a^{\prime}\bar{o}}^{\prime\prime}=y_{a^{\prime}\bar{o}}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and all $\bar{o}\in\mathcal{O}$. Since $q_{\varnothing}\geq\left|\mathcal{A}\right|$, $y^{\prime\prime}$ is an assignment. Moreover, since $R\left(o,r_{a}\right)\geq R\left(\varnothing,r_{a}\right)$,
$RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)\geq RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)\text{.}$
Then,
$RV\left(y,\left(r_{a},p_{-a}\right)\right)>RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)$
contradicts $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. Q.E.D.
Claim 2
Suppose that there is $y^{\ast}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}^{\ast}=1$ and $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$. Then, $\left|\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\right|\leq\left|\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)\right|$.
Proof. Suppose that there is $y^{\ast}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}^{\ast}=1$ and $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$. Then, by Claim 1, $y^{\ast}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
Next, we show that if $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$, then either $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ or $y^{\prime}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right),$ such that $y_{a\varnothing}^{\prime}=1$ and $y_{a^{\prime}o}^{\prime}=y_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$. Since $q_{\varnothing}\geq\left|\mathcal{A}\right|$, $y^{\prime}\in\mathcal{Y}$.
Suppose that $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$ and $y\notin\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. We have
$RV\left(y,\left(d_{a},p_{-a}\right)\right)=RV\left(y^{\ast},\left(d_{a},p_{-a}\right)\right)=RV\left(y^{\ast},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right).$
Then, by the construction of $d_{a}$, $y_{ao^{\prime}}=1\,$such that $R\left(o^{\prime},r_{a}\right)\geq R\left(\varnothing,r_{a}\right)$ and $R\left(o^{\prime},d_{a}\right)\geq R\left(\varnothing,r_{a}\right)$. Then, we have
$RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)\leq RV\left(y,\left(d_{a},p_{-a}\right)\right)\leq RV\left(y^{\ast},\left(d_{a},p_{-a}\right)\right)\text{.}$
Since $y^{\ast}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$, $y^{\prime}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. Therefore, if $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$, then either $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ or $y^{\prime}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ defined above.
Finally, we assume $\left|\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\right|\geq 2$ and arbitrarily let $y_{1},y_{2}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$ such that $y_{1}\neq y_{2}$. For $i=1,2$, let $y_{i}^{\prime}$ be such that $\left(y_{i}^{\prime}\right)_{a\varnothing}=1$ and $\left(y_{i}^{\prime}\right)_{a^{\prime}o}=\left(y_{i}\right)_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$.
We show $y_{1}^{\prime}\neq y_{2}^{\prime}$. Suppose not; that is, $y_{1}^{\prime}=y_{2}^{\prime}$. Then, $\left(y_{1}\right)_{a^{\prime}o}=\left(y_{2}\right)_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$. Since $y_{1}\neq y_{2}$, we have $o_{1}\neq o_{2}$ where $\left(y_{1}\right)_{ao_{1}}=1$ and $\left(y_{2}\right)_{ao_{2}}=1$. Without loss of generality, we assume $R\left(o_{1},d_{a}\right)<R\left(o_{2},d_{a}\right)$. However, $R\left(o_{1},d_{a}\right)<R\left(o_{2},d_{a}\right)$ and $\left(y_{1}\right)_{a^{\prime}o}=\left(y_{2}\right)_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$ imply
$RV\left(y_{1},\left(d_{a},p_{-a}\right)\right)<RV\left(y_{2},\left(d_{a},p_{-a}\right)\right)$
contradicting that $y_{2}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$. Therefore, $y_{1}^{\prime}\neq y_{2}^{\prime}$.
By these results, if $\left\{y_{1},y_{2},\cdots,y_{n}\right\}=\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right),$ then $\left\{y_{1}^{\prime\prime},y_{2}^{\prime\prime},\cdots,y_{n}^{\prime\prime}\right\}\subseteq\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ where $y_{i}^{\prime\prime}$ is either $y_{i}$ or $y_{i}^{\prime}$ constructed above. Since $y_{1}^{\prime},y_{2}^{\prime},\cdots,y_{n}^{\prime}$ are distinct, $y_{1}^{\prime\prime},y_{2}^{\prime\prime},\cdots,y_{n}^{\prime\prime}$ are also distinct and thus $\left|\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\right|\leq\left|\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)\right|$. Q.E.D.
Now, we show that $a$ weakly prefers $g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$. By Claim 2,
$\left|\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\right|\leq\left|\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)\right|\text{.}$
Moreover, by Claim 1, any $y^{\ast}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}^{\ast}=1$ with $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$ is also an element of $\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
Thus, for any $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$,
$\left(g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)\right)_{ao}$ $=$ $\left(f^{U}\left(d_{a},p_{-a}\right)\right)_{ao}\geq$
$\left(f^{U}\left(r_{a},p_{-a}\right)\right)_{ao}$ $=$ $\left(g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)\right)_{ao}\text{.}$
Since $g\left(f^{U}\left(d_{a},p_{-a}\right)\right)_{ao^{\prime}}=0$ for all $R\left(o^{\prime},r_{a}\right)>R\left(\varnothing,r_{a}\right)$,
$\left(g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)\right)_{a\varnothing}\leq\left(g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)\right)_{a\varnothing}.$
Thus, $a$ weakly prefers $g\left(f\left(d_{a},p_{-a}\right),r\right)$ to $g\left(f\left(r_{a},p_{-a}\right),r\right)$.
Proof of Theorem 2
[…]
First, we show that for all $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$, $y_{a\varnothing}=0$. Suppose not; that is, there is $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$, $y_{a\hat{o}}=1$ where $R\left(\hat{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)$. Then, by (3), there is $a^{\prime}\in\mathcal{A}$ such that $y_{a^{\prime}o^{\prime\prime}}=1$ with $R\left(o^{\prime\prime},p_{a^{\prime}}\right)>R\left(o^{\prime},p_{a^{\prime}}\right)\left(=R\left(\varnothing,r_{a}\right)\right),$ because $p_{a^{\prime}}=d_{a}$ for all $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$. Now, let $y^{\prime}\in\mathcal{Y}$ such that $y_{a^{\prime}\hat{o}}^{\prime}=1,$ $y_{a\varnothing}^{\prime}=1,$ and $y_{a^{\prime\prime}\bar{o}}^{\prime}=y_{a^{\prime\prime}\bar{o}}$ for all $a^{\prime\prime}\in\mathcal{A\setminus}\left\{a,a^{\prime}\right\}$ and all $\bar{o}\in\mathcal{O}$. Then, since
$R\left(\hat{o},r_{a^{\prime}}\right)=R\left(\hat{o},p_{a^{\prime}}\right)\text{ and }R\left(o^{\prime\prime},p_{a^{\prime}}\right)>R\left(\varnothing,r_{a}\right),$
$RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right),$
contradicting $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. Therefore, for all $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$, $y_{a\varnothing}=0$.
By this result,
$\left(f^{U}\left(r_{a},p_{-a}\right),r\right)_{a\varnothing}=\left(g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)\right)_{a\varnothing}=1$
On the other hand, since $f^{U}$ satisfies ETE,
$\left(f^{U}\left(d_{a},p_{-a}\right),r\right)_{a\hat{o}}=\left(g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)\right)_{a\hat{o}}>0.$
Therefore, $a$ strictly prefers $g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$.
Proof of Proposition 3
Suppose that there is $a\in\mathcal{A}$ whose true preference $r_{a}$ satisfies
$R\left(o,r_{a}\right)$ $<$ $R\left(\varnothing,r_{a}\right)<R\left(o^{\prime},r_{a}\right)$
$\sum\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}$$\geq$ $\left|\mathcal{A}\right|,$ (9)
for some $o,o^{\prime}\in\mathcal{O}$. Fix $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ arbitrarily. Moreover, we let $d_{a}\in\mathcal{D}_{a}$ be such that $R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},d_{a}\right)$.
We show the following fact.
Claim 3
Let $y$ and $y^{\prime}$ be such that $y_{a\varnothing}=y_{ao^{\prime}}^{\prime}=1$ and $y_{a^{\prime}\bar{o}}=y_{a^{\prime}\bar{o}}^{\prime}$ for all $a^{\prime}\in\mathcal{A}$ and all $\bar{o}\in\mathcal{O}$. Then, $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ if and only if $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
Proof. By the construction,
$RV\left(y,\left(r_{a},p_{-a}\right)\right)=RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)\text{.}$
First, suppose $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. Then, by (9), $y^{\prime}$ is an allocation. Suppose not; that is, there is $y^{\prime\prime}\in\mathcal{Y}$ such that $RV\left(y^{\prime\prime},\left(d_{a},p_{-a}\right)\right)<RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)$. Then, by (9), $o^{\prime\prime}$ such that $y_{ao^{\prime\prime}}^{\prime\prime}=1$ satisfies $R\left(o,d_{a}\right)\leq R\left(o^{\prime},d_{a}\right)$. Let
$o^{\prime\prime\prime}=o^{\prime\prime}\text{ if }o\neq o^{\prime},\text{ }o^{\prime\prime\prime}=\varnothing\text{ if }o^{\prime\prime}=o^{\prime}\text{.}$
Moreover, let $y^{\prime\prime\prime}$ be such that $y_{ao^{\prime\prime\prime}}^{\prime\prime\prime}=1$ and $y_{a^{\prime}\bar{o}}^{\prime\prime\prime}=y_{a^{\prime}\bar{o}}^{\prime\prime}$ for all $a^{\prime}\in\mathcal{A}$ and all $\bar{o}\in\mathcal{O}$. Then,
$RV\left(y^{\prime\prime\prime},\left(r_{a},p_{-a}\right)\right)\leq RV\left(y^{\prime\prime},\left(d_{a},p_{-a}\right)\right)<RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)=RV\left(y,\left(r_{a},p_{-a}\right)\right)$
contradicting $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$.
Second, suppose $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$. Suppose not; that is, there is $y^{\prime\prime}\in\mathcal{Y}$ such that $RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right)$. Then, by (9), $o^{\prime\prime}\in\mathcal{O}$ such that $y_{ao^{\prime\prime}}^{\prime\prime}=1$ and $R\left(o^{\prime\prime},r_{a}\right)\leq R\left(\varnothing,r_{a}\right)$. Let
$o^{\prime\prime\prime}=o^{\prime\prime}\text{ if }o^{\prime\prime}\neq\varnothing,\text{ }o^{\prime\prime\prime}=\bar{o}\text{ if }o^{\prime\prime}=\varnothing\text{,}$
where $\bar{o}$ satisfies $\sum y_{a\bar{o}}^{\prime\prime}<q_{\bar{o}}$ and $R\left(\bar{o},d_{a}\right)\leq R\left(o^{\prime},d_{a}\right)$. There exist such an $\bar{o}$, because (9) is satisfied. Then, let $y^{\prime\prime\prime}$ be such that $y_{ao^{\prime\prime\prime}}^{\prime\prime\prime}=1$ and $y_{a^{\prime}\bar{o}}^{\prime\prime\prime}=y_{a^{\prime}\bar{o}}^{\prime\prime}$ for all $a^{\prime}\in\mathcal{A}$ and all $\bar{o}\in\mathcal{O}$. Then,
$RV\left(y^{\prime\prime\prime},\left(d_{a},p_{-a}\right)\right)\leq RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)=RV\left(y,\left(d_{a},p_{-a}\right)\right)$
contradicting $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
Therefore, $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ if and only if $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$. Q.E.D.
By Claims 1 and 3, we have Proposition 3.
Proof of Lemma 4
We assume that the assumptions of Lemma 4 is satisfied. First, by Proposition 3, for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and all $\bar{o}\in\mathcal{O}$,
$g\left(f^{U}\left(e_{a},p_{-a}\right),r\right)_{a\bar{o}}=g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)_{a\bar{o}}.$
Second, by Theorem 2, $r_{a}$ is strategically dominated by $d_{a}$ with the refusal option. Therefore, $e_{a}$ is also strategically dominated by $d_{a}$ with the refusal option.
Proof of Proposition 4
We use the example introduced in the proof of Theorem 2. Suppose that there are $a\in\mathcal{A}$, $o,o^{\prime}\in\mathcal{O}$ satisfying (3). Let $d_{a}$ be such that $R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},d_{a}\right)$ and $p_{a^{\prime}}=d_{a}$ for all $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$. Then, since (3) is satisfied and $f^{U}$ satisfies ETE, there is $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$, $y_{ao^{\prime}}=1$ and $y_{a^{\prime}o^{\prime\prime}}=1$ for some $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$ and $o^{\prime\prime}\in\mathcal{O}$ such that $R\left(o^{\prime},d_{a}\right)<R\left(o^{\prime\prime},d_{a}\right)$. Then, if $y$ is realized, $a$ refuses $o^{\prime}$ but $a^{\prime}$ prefers $o^{\prime}$ to $o^{\prime\prime}$. Therefore, $g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)$ is wasteful.
Proof of Proposition 5
First, suppose that $p$ satisfies all of Conditions (i), (ii) and (iii). Then, for any two agents $a_{1}$ and $a_{2}$ belonging to different groups, $p_{a_{1}}$ and $p_{a_{2}}$ are not essentially the same. Moreover, for any two agents $a_{1}$ and $a_{2}$ belonging to the same group (either the second group or the third group), $\left(f^{M}\left(p\right)\right)_{a_{1}o^{\prime}}=\left(f^{M}\left(p\right)\right)_{a_{2}o^{\prime}}$ for all $o^{\prime}\in\mathcal{O}$.
Second if $p$ does not satisfies all of Conditions (i), (ii) and (iii), then $f^{M}\left(p\right)=f^{U}\left(p\right)$. By Proposition 2, $f^{M}$ satisfies ETE.
Proof of Proposition 6
First, if $r_{a}^{1}=\varnothing,$ then $\left(f^{M}\left(r_{a},p_{-a}\right)\right)_{a\varnothing}=1$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$. Thus, in this case, there is no $p_{a}\in\mathcal{P}$ that strategically dominate $r_{a}$. Thus, we assume $r_{a}^{1}\neq\varnothing$. We consider $p_{a}\left(\neq r_{a}\right)$. Let $k$ be the smallest integer such that $p_{a}^{k}\neq r_{a}^{k}$.
First, suppose $k\leq\bar{k}\left(p_{a}\right)$ and $r_{a}^{k}\neq\varnothing$. Since $k\leq\bar{k}\left(p_{a}\right)$, $r_{a}^{k^{\prime}}\neq\varnothing$ for all $k^{\prime}=1,\cdots,k-1$; that is, $R(\varnothing,r_{a})>k$ and $R(\varnothing,p_{a})\geq k$. In this case, since $f^{M}$ satisfies ETE (by Proposition 5), we revisit the examples in the first result of Lemma 3. When $k=1$ or $2$, trivially, we can use the example as is, because
$\left(g\left(f^{M}\left(r_{a},p_{-a}\right),r\right)\right)_{ao}$ $=$ $\left(f^{M}\left(r_{a},p_{-a}\right),r\right)_{ao},$
$\left(g\left(f^{M}\left(p_{a},p_{-a}\right),r\right)\right)_{ao}$ $=$ $\left(f^{M}\left(p_{a},p_{-a}\right),r\right)_{ao}.$
for all $o\in\mathcal{O}$.
We consider $k=3$. Then, since $R(\varnothing,r_{a})>k$ and $R(\varnothing,p_{a})\geq k$,
$\left(g\left(f^{M}\left(r_{a},p_{-a}\right),r\right)\right)_{ar_{a}^{1}}$ $=$ $\frac{q_{r_{a}^{1}}}{q_{r_{a}^{1}}+1},\text{ }\left(g\left(f^{M}\left(r_{a},p_{-a}\right),r\right)\right)_{ar_{a}^{k}}=\frac{1}{q_{r_{a}^{1}}+1}.$
$\left(g\left(f^{M}\left(r_{a},p_{-a}\right),r\right)\right)_{ap_{a}^{1}}$ $=$ $\frac{q_{r_{a}^{1}}}{q_{r_{a}^{1}}+1},$
and moreover, either
$\left(g\left(f^{M}\left(p_{a},p_{-a}\right),r\right)\right)_{ap_{a}^{k}}$ $=$ $\frac{1}{q_{r_{a}^{1}}+1}\text{ or}$
$\left(g\left(f^{M}\left(p_{a},p_{-a}\right),r\right)\right)_{a\varnothing}$ $=$ $\frac{1}{q_{r_{a}^{1}}+1}\text{.}$
Since $R(\varnothing,r_{a})>k$ and $R(p_{a}^{k},r_{a})>k$, in either case, $a$ prefers $g\left(f^{M}\left(r_{a},p_{-a}\right),r\right)$ to $g\left(f^{M}\left(p_{a},p_{-a}\right),r\right)$.
Second, suppose $k\leq\bar{k}\left(p_{a}\right)$ and $r_{a}^{k}=\varnothing$. Then, since we assume $r_{a}^{1}\neq\varnothing$, $k\geq 2$. Moreover, in this case, $p_{a}^{k^{\prime}}\neq\varnothing$ for all $k^{\prime}=1,\cdots,k$. Therefore, $R(\varnothing,p_{a})>R(\varnothing,r_{a})=k$. Let $p_{-a}$ be such that there are $q_{o}$ students represented by $a^{\prime}$ whose revealed preference is equivalent to $r_{a}(=p_{a^{\prime}})$ and the first best type of the other agents represented by $a^{\prime\prime}$ is $\varnothing$. Then, $\left(p_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii) with $l=k$, because $k\leq\bar{k}\left(p_{a}\right)$. Moreover, $\left(r_{a},p_{-a}\right)$ does not satisfy all of them. Therefore, in this case, $g\left(f^{M}\left(r_{a},p_{-a}\right)\right)$ is more preferable for $a$ than $g\left(f^{M}\left(p_{a},p_{-a}\right)\right)$.
Third, suppose $k>\bar{k}\left(p_{a}\right)$. In this case, $r_{a}$ and $p_{a}$ are essentially the same. Since $f^{M}$ satisfies ETE (by Proposition 5),
$g\left(f^{M}\left(r_{a},p_{-a}\right),r\right)=g\left(f^{M}\left(p_{a},p_{-a}\right),r\right)$
for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$.
Therefore, for any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ that strategically dominate $r_{a}$.